# Setup Once, Secure Always: A Single-Setup Secure Federated Learning Aggregation Protocol with Forward and Backward Secrecy for Dynamic Users

### v3

@@ Abstract.
In this paper, we propose a new secure aggregation protocol that requires only one setup operation for the entire FL training and allows new users to join or leave at any round. It employs lightweight symmetric homomorphic encryption with a key negation technique to efficiently mask updates, without user-to-user communication— unlike the existing protocols. To defend against model inconsistency attacks, we introduce a simple verification mechanism using message authentication codes (MACs). Our protocol is the first to combine forward/backward secrecy, dropout resilience, and model integrity verification in a single-setup design. We provide formal security proofs and implement an end-to-end prototype, which source code has been released [^1] Our source code is available at https://anonymous.4open.science/r/scatesfl-44E1 for performance testing, and at https://anonymous.4open.science/r/scatesfl-815A for real-world accuracy evaluation.. Our experimental results show that our protocol reduces user-side computation by approximately $99\%$ compared to state-of-the-art protocols like e-SeaFL (ACSAC’24), making it highly practical for real-world FL deployments, especially on resource-constrained devices.
@@ 8.2.1. Accuracy Comparison
Figures 8(a) and 8(b) present the accuracy of our protocol during training compared to the baseline FedAvg algorithm (McMahan et al., 2017), over 500 rounds with 100 users. While FedAvg achieves slightly higher final accuracy, our protocol demonstrates steady and reliable learning throughout the process. On CIFAR-10, our protocol reaches 38% accuracy versus FedAvg’s 45%, and on MNIST it attains 89.9%, which is close to FedAvg’s 96%. The small drop in accuracy comes from encoding floating-point values as integers, where a minor loss of precision accumulates across rounds. Despite this, our protocol maintains strong accuracy while providing additional benefits in privacy, communication efficiency, and reliability, making it a practical and effective choice in settings with system constraints or strict security requirements.
@@ Appendix A Detailed Analysis of Forward and Backward Secrecy in e-SeaFL and Flamingo
As discussed earlier, single-setup secure aggregation protocols offer a promising direction for scalable, efficient, and secure FL. Compared to multi-setup protocols, they significantly reduce communication and computation overhead by avoiding repeated cryptographic setup in each training round. Two notable protocols in this category are e-SeaFL (Behnia et al., 2024) and Flamingo (Ma et al., 2023), both of which adopt similar masking techniques based on long-term shared secrets and PRFs.
Among the two, e-SeaFL demonstrates improved performance and practicality due to its architectural design, which eliminates direct user-to-user interactions and relies instead on a few assisting nodes for mask coordination and recovery. This makes it more suitable for large-scale FL deployments.
Despite their efficiency, both e-SeaFL and Flamingo share a critical limitation: they do not provide FS and BS. In the following section, we analyze e-SeaFL in detail to demonstrate how its use of long-term cryptographic material exposes it to FS and BS vulnerabilities. Since Flamingo employs a similar masking strategy, the same analysis applies, indicating that Flamingo also lacks both FS and BS.
@@ A.1. Mask Derivation in e-SeaFL
e-SeaFL is a single-setup secure aggregation protocol proposed by Behnia et al. (ACSAC 2024) to improve communication efficiency by avoiding repeated cryptographic setup across rounds. In the setup phase, each user $i$ and assisting node $a\in\mathcal{A}_{i}$ perform a key agreement protocol ($\mathsf{KA}$) to derive a long-term shared secret or seed:
@@
where $SK_{i}$ and $PK_{a}$ are the private key of user $i$ and the public key of assisting node $a$, respectively. This shared secret $s_{i,a}$ is then used in each training round $t$ as input to a $\mathtt{PRF}$ to generate a round-specific mask value:
@@
After receiving all the aggregated masking contributions from the assisting nodes, the central server removes these masks from the aggregated global model parameters to recover the plaintext model update.
@@ A.2. Forward and Backward Secrecy Limitations
In e-SeaFL, round-specific masking values $m_{i,a}^{(t)}$ are deterministically derived from a long-term shared secret $s_{i,a}$ established between user $i$ and assisting node $a$. The masking function is defined in the Eq. 9. This design eliminates the need for repeated setup phases, but it introduces a critical vulnerability: if the shared secret $s_{i,a}$ is compromised, an adversary can compute all past (and future as well) masking values for every round. Consequently, the attacker can reconstruct the original model update $x_{i}^{(t)}$ for any round $t$ by subtracting the derived masks from the masked update $\tilde{x}_{i}^{(t)}$:
$x_{i}^{(t)}=\tilde{x}_{i}^{(t)}-\sum_{a\in\mathcal{A}_{i}}m_{i,a}^{(t)},$
where $\mathcal{A}_{i}$ is the set of assisting nodes for user $i$.
@@ A.3. Practical Implications
The absence of forward and backward secrecy in e-SeaFL and Flamingo present a tangible risk in real-world FL deployments. Devices such as smartphones, edge nodes, and IoT sensors often operate in environments with limited physical and software security, making them vulnerable to malware, unauthorized access, and insider threats- especially over prolonged training sessions.
Once a long-term secret $s_{i,a}$ is established between a user and an assisting node, it remains unchanged throughout the training process. If compromised at any point, an adversary can retroactively and proactively derive all masking values, enabling reconstruction of the user’s full update history. This undermines both privacy and the integrity of the aggregation process. While e-SeaFL and Flamingo offer efficiency through its single-setup design, its reliance on static secrets could make it unsuitable for deployments requiring strong long-term privacy guarantees.

### v4

@@ Abstract.
In this paper, we present a novel secure aggregation protocol that requires only a single setup for the entire FL training. Our protocol supports dynamic user participation, tolerates dropouts, and achieves both forward and backward secrecy. It leverages lightweight symmetric homomorphic encryption with a key negation technique to mask updates efficiently, eliminating the need for user-to-user communication. To defend against model inconsistency attacks, we introduce a low-overhead verification mechanism using message authentication codes (MACs). We provide formal security proofs under both semi-honest and malicious adversarial models and implement a full prototype. Experimental results show that our protocol reduces user-side computation by up to $99\%$ compared to state-of-the-art protocols like e-SeaFL (ACSAC’24), while maintaining competitive model accuracy. These features make our protocol highly practical for real-world FL deployments, especially on resource-constrained devices.
@@ 8.2.1. Accuracy Comparison
Figures 8(a) and 8(b) present the accuracy of our protocol during training compared to the baseline FedAvg algorithm (McMahan et al., 2017), over 500 rounds with 100 users. While FedAvg achieves slightly higher final accuracy, our protocol demonstrates steady and reliable learning throughout the process. On CIFAR-10, our protocol reaches 38% accuracy versus FedAvg’s 45%, and on MNIST it attains 89.9%, which is close to FedAvg’s 96%. The small drop in accuracy comes from encoding floating-point values as integers, where a minor loss of precision accumulates across rounds. This trade-off can be mitigated through techniques such as fixed-point encoding(Mohassel and Zhang, 2017; Knott et al., 2021) or quantization-aware training (QAT) (Jacob et al., 2018; Gholami et al., 2021), which preserve numerical precision while remaining compatible with secure aggregation. Despite this, our protocol maintains strong accuracy while providing additional benefits in privacy, communication efficiency, and reliability, making it a practical and effective choice in settings with system constraints or strict security requirements.
@@ Appendix A Detailed Analysis of Forward and Backward Secrecy in e-SeaFL and Flamingo
As discussed earlier, single-setup secure aggregation protocols offer a promising direction for scalable, efficient, and secure FL. Compared to multi-setup protocols, they significantly reduce communication and computation overhead by avoiding repeated cryptographic setup in each training round. Two notable protocols in this category are e-SeaFL (Behnia et al., 2024) and Flamingo (Ma et al., 2023), both of which adopt similar masking techniques based on long-term shared secrets and PRFs. Among the two, e-SeaFL demonstrates improved performance and practicality due to its architectural design, which eliminates direct user-to-user interactions and relies instead on a few assisting nodes for mask coordination and recovery. This makes it more suitable for large-scale FL deployments.
Despite their efficiency, both e-SeaFL and Flamingo share a privacy concern: they do not provide FS and BS under their current single-setup design. In the following section, we analyze e-SeaFL in detail to demonstrate how its use of long-term cryptographic material exposes it to FS and BS vulnerabilities. Since Flamingo employs a similar masking strategy, the same analysis applies, indicating that Flamingo also lacks both FS and BS.
@@ A.1. Mask Derivation in e-SeaFL
In e-SeaFL, each user $i$ and assisting node $a\in\mathcal{A}_{i}$ perform a key agreement protocol ($\mathsf{KA}$) to derive a long-term shared secret or seed in the setup phase:
@@
where $SK_{i}$ and $PK_{a}$ are the private key of user $i$ and the public key of assisting node $a$, respectively and $\mathcal{A}_{i}$ is the set of participating assisting nodes. This shared secret $s_{i,a}$ is then used in each training round $t$ as input to a $\mathtt{PRF}$ to generate a round-specific mask value:
@@
After receiving all the aggregated masking contributions from the assisting nodes, the central server removes these masks from the aggregated global model parameters to recover the plaintext model update, as follows:
$\theta=\sum_{i\in\mathcal{U}}x_{i}^{(t)}=\sum_{i\in\mathcal{U}}\tilde{x}_{i}^{(t)}-\sum_{a\in\mathcal{A}_{i}}\left(\sum_{i\in\mathcal{U}}m_{i,a}^{(t)}\right)$
@@ A.2. Forward and Backward Secrecy Limitations
In e-SeaFL, round-specific masking values $m_{i,a}^{(t)}$ are deterministically derived from long-term shared secrets $s_{i,a}$ between user $i$ and assisting node $a$, as defined in Eq. 9. If only a subset of these secrets is compromised, i.e. $\mathcal{A}^{\prime}_{i}\subset\mathcal{A}_{i}$, the adversary can compute only partial masking shares. Since at least one share remains unknown, the masked update retains information-theoretic secrecy. This threshold-like structure provides resilience against partial compromise and is a notable strength of e-SeaFL.
However, this protection collapses entirely if all long-term secrets $s_{i,a}$ for $a\in\mathcal{A}_{i}$ are compromised. The adversary can retroactively and proactively compute all masking shares $m_{i,a}^{(t)}$ across every round $t$, enabling full reconstruction:
(10) $x_{i}^{(t)}=\tilde{x}_{i}^{(t)}-\sum_{a\in\mathcal{A}_{i}}m_{i,a}^{(t)}.$
In other words:
  • Forward secrecy is broken: Compromising the long-term secrets at time $T$ enables recovery of all past updates ($t<T$).
  • Backward secrecy is broken: The same compromise allows prediction of all future masks ($t>T$) and recovery of future updates.
@@ A.3. Practical Implications
The absence of FS and BS in e-SeaFL (similarly in Flamingo) introduces a critical privacy concern in real-world FL deployments. Devices such as smartphones, edge nodes, and IoT sensors frequently operate in hostile environments— exposed to malware, physical tampering, and insider threats. Over prolonged training sessions, the probability of compromise increases significantly.
Once a long-term secret $s_{i,a}$ is established between a user and an assisting node, it remains static throughout the entire training lifecycle. If compromised, even once, an adversary can deterministically regenerate all masking values $m_{i,a}^{(t)}$ across every round. This enables full reconstruction of the user’s masked updates. The breach is not limited to a single round; it retroactively exposes past contributions and proactively jeopardizes future ones. While e-SeaFL achieve efficiency through single-setup designs, its reliance on static secrets makes it unsuitable for deployments demanding robust confidentiality.
In contrast, our protocol eliminates this risk entirely. By using fresh, per-round randomness and avoiding long-term key dependencies, we ensure that each training round is cryptographically isolated. Even if a device is compromised, the damage is strictly contained to that round. There is no possibility of past and future exposures. This design offers not just theoretical security, but practical resilience without requiring complex key ratcheting, proactive key refreshes, or threshold-based updates.